OpenStudy (anonymous):

how to find the x and y intercepts of y=x-5

OpenStudy (anonymous):

Set x and y equal to 0 separately. y = x - 5 y = 0 - 5 Solve here and this will be the y-intercept. y = x - 5 0 = x - 5 Solve here and this will be the x-intercept.

OpenStudy (anonymous):

@Calmathlete so the x intercept is ........ the y intercept is -5

OpenStudy (anonymous):

The y-intercept is -5. Solve for x to get the x-intercept.

OpenStudy (anonymous):

x is 0?

OpenStudy (anonymous):

Nope... Get x by itself. What would you do to get rid of that -5?

OpenStudy (anonymous):

add 5 to both sides?

OpenStudy (anonymous):

Yup. If you do that, the -5 will cancel and x will equal...

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

i don't understand could you explain it please please?

jimthompson5910 (jim_thompson5910):

If you take 0 = x - 5 and add 5 to both sides, what do you get?

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

0 = x - 5 +5 +5

OpenStudy (anonymous):

@jim_thompson5910 we got 5 as x and to find y you just substitute x and then simplify?

jimthompson5910 (jim_thompson5910):

You are correct in saying that x = 5 is the x-intercept You can write the x-intercept as (5,0)

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean by "to find y you just substitute x and then simplify"

jimthompson5910 (jim_thompson5910):

you mean the y-intercept?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

To find the y-intercept, you plug in x = 0 and simplify

jimthompson5910 (jim_thompson5910):

y=x-5 y=0-5 ... replace x with 0 and simplify

OpenStudy (anonymous):

so y is -5?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thankyou so much yu cleared everything up :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910 one more question sorry how can u graph a ine using those intercepts? thankyou so much:)

jimthompson5910 (jim_thompson5910):

Above you found the x and y intercepts to be (5,0) and (0,-5) agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

These are two points that lie on the line. So plot these two points and draw a line through them to graph y = x-5